# Can LLMs Generate Diverse Molecules? Towards Alignment with Structural Diversity

### v1

@@ 4 Experiment
In this section, we validate our supervised fine-tuning and reinforcement learning methods for generating diverse molecules, coined Div-SFT and Div-SFT+RL, respectively. In these experiments, one can observe that (1) our fine-tuning approach enables LLMs to better discover diverse molecules compared to decoding schemes for diversified generation and (2) our fine-tuned LLM outperforms other representative LLMs, including generalist and specialist models for chemical tasks.
@@ 4.1 Comparison with decoding schemes
In this experiment, we show that our method enables LLMs to better generate diverse molecules compared to the existing decoding schemes for diverse sequence generation. To validate the consistent improvement, we implement our fine-tuning method and decoding schemes on two models specialized in the chemical domain: BioT5+ (Pei et al., 2024) and MolT5 (Edwards et al., 2022).
(a) Comparison with decoding schemes on BioT5+
(b) Comparison with decoding schemes on MolT5
Figure 5: Comparison with decoding schemes. NCircles represents both quality and diversity-related metric. Top 10 and Accepted & unique represent quality-related metrics. IntDiv. represents a diversity-related metric. Our method generates more diverse and high-quality molecules compared to the baselines. Notably, our method makes a larger gap over the baselines on metrics related to capturing both quality and diversity, i.e., NCircles.
Tasks. We consider description-guided molecule generation using the ChEBI-20 dataset, which includes training and test sets (Edwards et al., 2021). Note that the training dataset has also been used to pre-train the base LLMs, BioT5+ and MolT5, i.e., not an external dataset. The test dataset is unobserved during both the pre-training and fine-tuning phases. Each data point consists of a pair of a description and a ground-truth example molecule that satisfies the description.
In our experiments, we consider generating $50$ molecules for each description in the ChEBI-20 test dataset involving $3300$ molecular descriptions. We then evaluate both the structural diversity of the generated molecules and how well they satisfy the given descriptions.
Metrics. To measure the similarity between two molecules, we use Tanimoto similarity (Bajusz et al., 2015), which based on distance of molecular structures from fingerprints. To be specific, we consider the following metrics for generated molecules:
  • The number of accepted and unique molecules (Accepted & unique): This metric measures the number of valid and unique molecules satisfying the given description. [^3] Following the evaluation metric of prior studies (Edwards et al., 2022; Pei et al., 2024), we evaluate the molecule as satisfying the given description when its BLEU score between a ground-truth sample is high ($>0.7$).
  • The number of circles (NCircles; Xie et al., 2023): This metric considers both the quality and diversity of generated molecules. For the accepted molecules satisfying the given description, this metric counts the number of maximum distinct clusters with Tanimoto similarity (Bajusz et al., 2015) below a threshold $h$, i.e., indicating chemical space coverage (illustrated in Figure 6).
@@
  • Internal diversity (IntDiv.; Polykovskiy et al., 2020): This metric is the complement of the average of pair-wise Tanimoto similarities between the accepted molecules satisfying the given description.
  • Average of top 10 scores (Top 10): This metric measures the quality of the generated molecules by averaging the BLEU scores (Papineni et al., 2002) of the top unique $10$ molecules yielding high BLEU scores between the ground-truth example.
(a) BS
(22,11)
(b) Diverse BS
(7,4)
(c) Contrastive BS
(26, 11)
(d) Div-SFT+RL
(29,28)
Figure 6: The distribution of (Accepted & unique, NCirclesh=0.65) obtained by each method. The molecules are generated from the description in Table 13. The Fruchterman-Reingold force-directed algorithm (Fruchterman & Reingold, 1991) embeds each molecule in a 2-dimensional space based on Tanimoto similarity. Here, a low distance yields a high Tanimoto similarity. Each dot, shaded circle, and edge represent accepted molecules, distinct clusters (used to measure NCircles), and a pair with similarity above $0.65$, respectively. Our method better captures the molecular diversity.
Table 1: Visualization of the generated molecules with their diversity. The eight molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.” The generated molecule with blue line indicates the accepted molecule.
Method (IntDiv.) Example of outputs
BS (0.51) [Uncaptioned image]
Diverse BS (0.28) [Uncaptioned image]
Contrastive BS (0.53) [Uncaptioned image]
Div-SFT+RL (0.69) [Uncaptioned image]
@@
Implementations. For supervised fine-tuning, we collect a hundred molecules for each molecular description $p_{\text{desc}}$ in ChEBI-20 training sets. Then, they are filtered to remove invalid molecules, duplicated molecules, and unaccepted molecules. The description-matching reward $r_{\text{match}}(m_{k})$ is defined as a BLEU score between a ground-truth example. The diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ is defined as the complement of the maximum Tanimoto similarity between generated molecules $\{m_{i}\}^{k-1}_{i=1}$. The detailed implementations and hyper-parameters follow Appendix A.
@@
Baselines. We compare our fine-tuning approach with various decoding schemes. We consider the random sampling with different temperatures $\{0.7,1.0,1.5\}$ and nucleus sampling (Holtzman et al., 2020). We also consider beam search (BS) and the variants of BS to promote sequence-level diversity between the generated samples: diverse BS (Vijayakumar et al., 2018) and contrastive BS (Su et al., 2022). The detailed settings are described in Appendix B.
Results. We present the main results on BioT5+ and MolT5 in Figure 5(a) and Figure 5(b), respectively. One can see that our approach, i.e., Div-SFT+RL, shows superior performance compared to the considered baselines. Especially, it is worth noting that our approach makes significant improvements in NCircles metrics that require generating diverse molecules satisfying the given description.
We also provide the qualitative results in Figure 6 and Table 13 with sampled molecules. One can see that variants of beam search for diverse sentence generation, i.e., diverse BS and contrastive BS, do not enhance molecular structural diversity. In contrast, our approach demonstrates significantly higher molecular structural diversity. In Appendix C, we provide additional examples of molecules generated by the considered baselines and our method.
Table 2: Comparison with existing LLMs on description-based molecule generation. Our fine-tuned model shows superior performance compared to the considered baselines.
Method Accepted & Unique NCirclesh=0.75 NCirclesh=0.65 IntDiv. Top 10
Chemical-task Specialist LLMs
MolT5 3.12 1.85 1.59 0.29 0.41
Text+Chem T5 13.78 4.82 3.29 0.23 0.63
BioT5 15.58 8.45 6.07 0.33 0.69
BioT5+ 16.03 7.93 6.16 0.31 0.63
Generalist LLMs
Mol-instructions 0.06 0.03 0.03 0.02 0.01
LlaSMol 17.68 6.73 4.94 0.26 0.68
GPT-3.5-turbo 3.87 2.83 2.49 0.20 0.41
GPT-4o 5.74 4.09 3.53 0.22 0.48
o1-preview 6.92 3.46 2.36 0.13 0.31
BioT5++ours 21.98 16.98 14.35 0.45 0.74
4.2 Comparison with existing LLMs
Here, we compare our fine-tuned BioT5+, presented in Figure 5, with various recent LLMs, including chemical-task specialists, generalists, and fine-tuned generalist LLMs on chemical domains. The purpose of this experiment is to highlight the limitations of existing LLMs in capturing molecular diversity, whereas our fine-tuned model successfully captures molecular diversity during generation. The tasks and metrics are the same as settings in Section 4.1. We use the first $500$ molecular descriptions in the ChEBI-20 test dataset for evaluation.
Baselines. We compare our fine-tuned model with various existing LLMs. We consider four LLMs specialized for chemical tasks: Text+Chem T5 (Christofidellis et al., 2023), MolT5, BioT5 (Pei et al., 2023), and BioT5+. Next, we consider two generalist LLMs: Mol-instructions (Fang et al., 2024) fine-tuned from Llama-7B (Touvron et al., 2023), and LlaSMol (Yu et al., 2024) fine-tuned from Mistral-7B (AI, 2023), based on description-based molecular generation tasks. For each baseline, we report the best results (highest NCirclesh=0.65) obtained using either random sampling, beam search, diverse beam search, or contrastive beam search. We also consider three strong API-based generalist LLMs: GPT-3.5, GPT-4o, and o1-preview (OpenAI, 2023; 2024). [^4] We use gpt-3.5-turbo-0125 and gpt-4o-2024-08-06. The detailed experimental settings and prompts are described in Appendix B.
Results. We present the results in Table 2. One can see that our fine-tuned model shows superior performance compared to the considered baselines in discovering diverse and high-quality molecules. Furthermore, it is noteworthy that most existing LLMs yield low NCircles with respect to the number of generated accepted and unique molecules, while our method yields relatively high NCircles.
4.3 Fine-tuning generalist LLMs
We further validate whether our fine-tuning method improves the generalist LLMs as well in terms of the diversity of generated molecules. Here, as a base LLM for implementing our method, we consider DrugAssist (Ye et al., 2023) which is fine-tuned from the Llama-7B. As baselines, we apply random sampling and contrastive beam search to DrugAssist.
Tasks. In this experiment, we consider prompts based on the four quantitative molecule properties: hydrogen bond (HB) donors, HB acceptors, Bertz complexity (Bertz, 1981), and quantitative estimate of drug-likeness (QED) (Bickerton et al., 2012). The goal of this task is to generate diverse molecules that satisfy the properties described in the given prompt.
Implementations. We apply our supervised fine-tuning and reinforcement learning to enhance DrugAssist. For supervised fine-tuning, we collect multiple molecules for prompts about three properties: HB donors, HB acceptors, and Bertz complexity. The prompts about QED are excluded from training but included in the evaluation to assess generalization to unseen properties. The sampled molecules are filtered to remove invalid molecules, duplicated molecules, and molecules that do not satisfy the given properties. The property-match reward $r_{\text{match}}(m_{k})$ yields a non-zero value when satisfying the given property, and the diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ is defined as same as Section 4.1. The detailed implementations and hyper-parameters follow Appendix A.
@@ 4.3 Fine-tuning generalist LLMs
Figure 7: Experiments on DrugAssist. Our method consistently improves the performance for generating diverse and high-quality molecules when implemented on generalist LLMs.
Results. We present the results in Figure 7. One can see that our approach consistently improves the performance in generating diverse and high-quality molecules when implemented on generalist LLMs. Note that our approach consistently demonstrates superior performance for the unseen prompt, i.e., the prompt about QED in Figure 7(d).
Table 3: Experiments with the large number of samples. Our method discovers more diverse molecules with respect to the (1) the number of generations and (2) time costs.
Methodnum. of ${}_{\text{num. of generations}}$ ${\text{BS}_{300}}$ ${\text{BS}_{400}}$ ${\text{BS}_{500}}$ ${\text{Ours}_{85}}$ ${\text{Ours}_{120}}$ ${\text{Ours}_{155}}$
NCirclesh=0.65 18.6 20.4 21.3 20.4 23.6 25.6
Time (sec.) 323 452 585 65 86 107
@@ 4.4 Ablation studies
Table 4: Comparison with variants of implementations. Applying multi-stage reinforcement learning shows superior performance compared to (1) supervised fine-tuning with hard filtering and (2) single-stage reinforcement learning.
Method Accepted & Unique NCirclesh=0.75 NCirclesh=0.65 IntDiv. Top 10

### v2

@@ 4 Experiment
In this section, we validate our supervised fine-tuning and reinforcement learning methods for diverse molecular generation, coined Div-SFT and Div-SFT+RL, respectively.
@@ 4.1 Comparison with Decoding Schemes
We first show that our fine-tuning method enables molecular generative LLMs to generate diverse molecules better than applying decoding schemes for diverse sequence generation. Here, we implement our method and decoding schemes on a recent description-guided molecular generative LLM: BioT5+ (Pei et al., 2024).
Tasks. We consider diverse molecular generation with two existing datasets: L+M-24 (Edwards et al., 2024) and ChEBI-20 (Edwards et al., 2021). Each data point in these datasets provides a qualitative molecular property description, e.g., “The molecule is an EGFR inhibitor”, and some examples of target molecules that satisfy the description. [^4] In L+M-24 training split, an average of $70$ provided molecules corresponds to each molecular property description. We provide the data statistics in Appendix C. Both L+M-24 and ChEBI-20 datasets consist of training and test splits where the training splits have been used to pre-train the base LLM. Our fine-tuning method also uses their training splits. We generate $50$ molecules for each description in the test splits for evaluation. [^5] In Section 4.4, we further validate our method by generating $100$, $150$, and $200$ molecules for each description.
Metrics. We evaluate the structural diversity of the generated molecules that satisfy the given molecular description. However, evaluating whether the generated molecule satisfies some qualitative properties, e.g., a membrane stabilizer, is intractable due to the complexity of the required assessments. Therefore, we assume that the molecule satisfies the description if it shares a certain degree of structures with one of the target molecules provided in the dataset. [^6] In Tables 2 and 4.3, we also evaluate whether the generated molecules actually satisfy the molecular description using an external tool, e.g., a docking tool. Here, the specific degree is described in the explanation of the accepted molecules.
To assess the similarity between two molecules, we compute Dice and Tanimoto similarities on their structural features (Bajusz et al., 2015), focusing on the degree of shared substructures and overall structural similarity, respectively. See Section A.2 for details. Based on these, we evaluate the following metrics for the generated molecules:
  • The number of accepted and unique molecules (Accepted & Unique) counts unique molecules that share substructures with one of the provided examples of target molecules (as much as Dice similarity higher than $0.7$).
  • The number of circles (NCircles; Xie et al., 2023) considers both quality and diversity. Given the set of accepted molecules, NCirclesh computes the size of the largest subset of molecules in which no two molecules are structurally similar to each other (as much as Tanimoto similarity below a threshold $h$).
@@
  • Internal diversity (IntDiv.; Polykovskiy et al., 2020) measures the average pair-wise structural distances, based on the complement of Tanimoto similarity, between the accepted molecules.
@@
Implementations. For supervised fine-tuning, we collect molecules for each description $p_{\text{desc}}$ in the training splits (a maximum of $100$ molecules per description). Then, the collected molecules are filtered to remove invalid, duplicated, and unaccepted molecules. The description-matching reward $r_{\text{match}}(m_{k})$ is defined as the maximum Dice similarity between the molecule $m_{k}$ and the provided target molecules. Next, the diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ is defined as the complement of the maximum Tanimoto similarity between the molecule $m_{k}$ and the previously generated molecules $\{m_{i}\}^{k-1}_{i=1}$. We describe detailed implementations in Appendix B.
@@
Baselines. We compare our method with various decoding schemes, including random sampling, nucleus sampling (Holtzman et al., 2020), and beam search (BS; Rosenberg and Baldwin, 1965). We also consider the variants of BS that promote sequence-level diversity: diverse BS (Vijayakumar et al., 2018) and contrastive BS (Su et al., 2022).
Results. In Table 1, we present experimental results on L+M-24 and ChEBI-20 datasets. One can see that applying our fine-tuning approach shows superior performance in discovering diverse accepted molecules, i.e., yields the highest NCircles, compared to applying existing decoding schemes on the molecular generative LLM. We also present qualitative results in Table 2, which evaluate whether the generated molecules actually satisfy the given description using an external tool (Trott and Olson, 2010). One can see that both BS and our method generate molecules that satisfy the given description, but our method generates more diverse molecules compared to BS.
Table 3: Comparison with results obtained by applying existing decoding schemes to each LLM. The base LLM of our method is BioT5+. For each baseline, we report the best result obtained by one of the existing decoding schemes. Our fine-tuned model shows superiority in generating diverse molecules by yielding highest NCircles.
Dataset Method NCirclesh=0.85 NCirclesh=0.75 Accepted & Unique IntDiv.
MolT5 (Diverse BS) 6.295 4.308 25.841 0.278
L+M-24 BioT5+ (Contrastive BS) 4.562 2.603 33.568 0.176
Meditron (Diverse BS) 4.693 3.037 27.837 0.263
Ours 10.511 6.278 31.062 0.287
MolT5 (Diverse BS) 1.902 1.509 2.798 0.101
Text+Chem T5 (Contrastive BS) 6.674 4.408 15.828 0.205
ChEBI-20 BioT5+ (Contrastive BS) 7.955 5.883 15.215 0.233
LlaSMol (BS) 7.371 5.492 16.978 0.217
Ours 11.301 8.966 16.271 0.246
@@
Figure 3: Experiments with fine-tuned generalist LLMs (DrugAssist; Ye et al., 2023). Our method consistently improves the performance for generating diverse and high-quality molecules.
Table 4: Experiments with the large number of samples. The base LLM is BioT5+. Our method discovers more diverse molecules with respect to the (1) the number of generations and (2) time costs.
Methodnum. of ${}_{\text{num. of generations}}$ ${\text{BS}_{300}}$ ${\text{BS}_{400}}$ ${\text{BS}_{500}}$ ${\text{Ours}_{100}}$ ${\text{Ours}_{150}}$ ${\text{Ours}_{200}}$
NCirclesh=0.85 13.971 14.553 15.821 14.472 17.215 19.134
Time (sec.) 323 452 585 75 107 146
4.2 Comparison with Existing LLMs
Additionally, we compare our fine-tuned BioT5+ in Section 4.1 with baselines that apply existing decoding schemes to other LLMs, extending the experiment in Section 4.1. The purpose of this experiment is to highlight that other existing LLMs have limitations in diverse molecular generation when relying on existing decoding schemes for diverse sequence generation, compared to our fine-tuned model. The overall metrics and tasks are the same as settings in Section 4.1.
Baselines. For the comparison on L+M-24 dataset, we additionally consider LLMs trained on L+M-24 training split: MolT5 (Edwards et al., 2024, 2022) and Meditron (Edwards et al., 2024; Chen et al., 2023). For the comparison on ChEBI-20 dataset, we additionally consider LLMs trained on ChEBI-20 training split: MolT5 (Edwards et al., 2022), Text+Chem T5 (Christofidellis et al., 2023), and LlaSMol (Yu et al., 2024). For each baseline, we report the best results (highest NCircles) obtained using one of existing decoding schemes, i.e., random sampling, BS, diverse BS, and constrastive BS. We describe detailed settings in Appendix D.
Results. We present the results in Table 3. One can see that most baselines yield a low NCircles with respect to the number of accepted and unique molecules, indicating a lack of diversity, while our method yields a relatively high NCircles.
Table 5: Comparison with variants of implementations. Our multi-stage reinforcement learning shows superior performance compared to the single-stage reinforcement learning.
Method NCirclesh=0.85 NCirclesh=0.75 Accepted & Unique IntDiv.
[…]
Div-SFT+RL (ours) 11.301 8.966 16.271 0.246
4.3 Fine-tuning Generalist LLMs
We further validate whether our fine-tuning method improves the generalist LLMs. Here, as a base LLM for implementing our method, we consider DrugAssist (Ye et al., 2023) which is based on the Llama-7B (Touvron et al., 2023). As baselines, we apply existing decoding schemes to DrugAssist.
Tasks. We consider molecular descriptions about quantitative molecular properties: hydrogen bond (HB) donors, HB acceptors, Bertz complexity (Bertz, 1981), and quantitative estimate of drug-likeness (QED) (Bickerton et al., 2012). Note that these properties can be evaluated using external tools like RDKit (Landrum et al., 2025).
Implementations. We consider fine-tuning the base LLM on prompts about three properties: HB donors, HB acceptors, and Bertz complexity. The prompts about QED are used to assess generalization to unseen properties. The overall implementations follow Section 4.1, but the description-matching reward $r_{\text{match}}(m_{k})$ is designed to yield a positive value when the given property is exactly satisfied. See Appendix B for the detailed settings.
Results. We present the results in Figure 3. One can see that our approach improves the performance in generating diverse molecules exactly satisfying the given molecular descriptions. Furthermore, our approach consistently demonstrates superior performance for the unseen prompt, i.e., the prompt about QED in Figure 3(d).
@@ 4.4 Ablation Studies
Large number of samples vs. performance. We also analyze how well our method discovers diverse molecules with respect to the number of generations. Here, we extend beyond the settings in Section 4.1. We consider our method to generate $100$, $150$, and $200$ molecules with a single NVIDIA A100 SXM4 40GB GPU. We also consider BS to generate a larger number of molecules with beam sizes of $300$, $400$, and $500$, reaching our maximum computational budget (four NVIDIA A100 SXM4 40GB GPUs). This experiment uses $250$ molecular descriptions in the ChEBI-20 test split. We present the results in Table 4. One can see that our method exhibits further performance improvements compared to the baseline, even though the baseline generates a larger number of molecules using our maximum computational budget.
Time costs vs. performance. In addition, we also analyze how well our method discovers diverse molecules with respect to the time costs. In Table 4, we present the time costs for each method. One can see that our method discovers more diverse molecules with respect to the time costs.
Table 6: Coverage of the target molecules in L+M-24 dataset. The base LLM is BioT5+.
Method Target Coverage
Random 0.137
BS 0.420
Contrastive BS 0.420
Diverse BS 0.343
Ours 0.558
Coverage of the target molecules. We further validate how our method benefits capturing the wide range of target molecular space. Here, we compute how many target molecules in the L+M-24 dataset (an average of $17$ molecules per description) can be captured within the generation space. The target molecule is considered captured if it is significantly similar to one of the generated molecules (Dice similarity $>0.95$). In Table 6, we present the average ratio of captured target molecules. One can see that our method yields the highest score.
5 Conclusion
In this paper, we identify the limitations of large language models (LLMs) for generating diverse molecules. In response, we present a new fine-tuning approach to adapt existing LLMs to generate diverse molecules. Experiments show that our approach enables LLMs to better discover diverse molecules compared to the existing approaches. This success highlights the potential of our method to advance LLM-driven drug discovery.
6 Limitations
Our method may require autoregressively generating a very long sequence to induce a huge set of molecules. In response, an interesting avenue for future work is to reduce the space and time complexity in generating the sequence of molecules, for example, introducing continuous tokens (Hao et al., 2025) that encode the set of previously generated molecules (Zaheer et al., 2017). Next, we validated whether the LLM-generated molecules satisfy the given molecular descriptions using computational approaches. However, this does not imply that the generated molecules will necessarily satisfy the given molecular descriptions in tricky real-world scenarios. To ensure the reliability of our method, the generated molecules should be validated through the real-world experiments. Additionally, although the molecules generated by each method are valid in computational terms, they may not be synthesizable in practice. Lastly, due to the limited computational budgets, we conducted experiments only on models with up to 7B parameters. The generalizability of our method to larger models (e.g., 70B) remains unexplored and is left for future work.
7 Ethical Considerations
Our fine-tuning method enables LLMs to generate diverse molecules from textual descriptions of molecular properties. However, these advancements also introduce potential risks, such as the generation of harmful drugs and the misuse of synthesized molecules.
Appendix A Additional Related Works
A.1 Reinforcement Learning (RL) for Diverse Molecular Generation
Existing literature has studied RL-based methods to generate molecules with desired properties while enhancing their diversity. First, Blaschke et al. (2020); Pereira et al. (2021) introduced memory-assisted RL, which penalizes the reward of a molecule when it is highly similar to the molecules stored in the memory unit. He et al. (2024) also incorporated RL with a diversity penalty in transformer-based architectures for molecular generation. In addition, Hu et al. (2024) leveraged multiple GPT-based agents trained with RL to encourage these agents to explore diverse directions for discovering diverse molecules. Their algorithms are designed to consider a fixed target property. In contrast, our work fine-tunes LLMs to generate diverse molecules given a prompt that is flexible to describe various target properties.
A.2 Molecular Similarity Measures
In this section, we explain the structural similarity between two molecules. These measures are used to define the reward of reinforcement learning (Appendix B) and diversity metrics (Section A.3).
Molecular structural features. First, molecular structural features are expressed with Morgan fingerprint (Rogers and Hahn, 2010), which is a vector that characterizes the presence of specific atoms, bonds, or substructures in a given molecule. Note that we get this from RDkit package (Landrum, 2016). We denote this with $f(m)$, which maps the molecule $m$ to its Morgan fingerprint. Next, we compute the molecular structural similarity based on the molecular fingerprints.
Tanimoto similarity. The overall similarity between two molecules is typically evaluated as follows:
$T(m_{i},m_{j})=\frac{|f(m_{i})\cap f(m_{i})|}{|f(m_{i})\cup f(m_{j})|},$
where $f(m_{i})$ maps the molecule $m_{i}$ to its Morgan fingerprint. This similarity $T(m_{i},m_{j})$ is referred to as Tanimoto similarity (Bajusz et al., 2015), evaluating overall structural similarity.
Dice similarity. In this paper, we also consider Dice similarity (Bajusz et al., 2015) which is sensitive to the degree of shared structural features between two molecules. This is defined as follows:
$D(m_{i},m_{j})=\frac{2|f(m_{i})\cap f(m_{i})|}{|f(m_{i})|+|f(m_{j})|},$